OpenStudy (r_josh10):

An item is priced at $13.77. If the sales tax is 4%, what does the item cost including sales tax? A. $19.28 B. $0.55 C. $28.09 D. $14.32

OpenStudy (r_josh10):

@Michele_Laino

OpenStudy (michele_laino):

by definition of percentage, we have to compute this: \[\large 13.77 + 13.77 \times \frac{4}{{100}} = ...?\]

OpenStudy (michele_laino):

since the sales tax is: 13.77*(4/100)

OpenStudy (r_josh10):

1.1016?

OpenStudy (michele_laino):

hint: what is: \[13.77 \times \frac{4}{{100}} = ...?\]

OpenStudy (r_josh10):

0.5508

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, what is: \[13.77 + 0.5508 = ...?\]

OpenStudy (r_josh10):

then what?

OpenStudy (r_josh10):

14.3208

OpenStudy (michele_laino):

that's right!